A comparative balance sheet and income statement is shown for Cruz, Inc.

CRUZ, INC.
Comparative Balance Sheets
December 31, 2017
2017 2016
Assets
Cash $ 85,600 $ 21,300
Accounts receivable, net 36,800 45,200
Inventory 77,100 84,900
Prepaid expenses 4,700 3,900
Total current assets 204,200 155,300
Furniture 94,700 110,500
Accum. depreciation—Furniture (14,700 ) (8,400 )
Total assets $ 284,200 $ 257,400
Liabilities and Equity
Accounts payable $ 13,400 $ 19,000
Wages payable 8,000 4,500
Income taxes payable 1,400 2,500
Total current liabilities 22,800 26,000
Notes payable (long-term) 28,900 66,400
Total liabilities 51,700 92,400
Equity
Common stock, $5 par value 204,000 162,300
Retained earnings 28,500 2,700
Total liabilities and equity $ 284,200 $ 257,400

   

CRUZ, INC.
Income Statement
For Year Ended December 31, 2017
Sales $ 440,700
Cost of goods sold 283,700
Gross profit 157,000
Operating expenses
Depreciation expense $ 33,900
Other expenses 80,400 114,300
Income before taxes 42,700
Income taxes expense 15,500
Net income $ 27,200


1. How much cash is paid to acquire inventory during year 2017?
  
Cash Paid for inventory:???

Use the above balance sheet and income statement to prepare the cash provided (used) from operating activities section by direct method. (Amounts to be deducted should be indicated with a minus sign.)

Cash flows from operating activities:
Answer #1

1) Cash Paid to Acquire Inventory(Purchased made during the year)

Opening Inventory + Purchase = Sales + Closing Inventory

84,900 + Purchase = 283,700 + 77,100

Purchase = 360,800 - 84,900

= 275,900 $

Cash Flow from Operating Activities
Cash Flow from Revenue (Net Receipts) 449,100
Less: Cash Payments for Expenses 273,700
Income Before Income Taxes 175,400
Less : Cash Payment for Income Taxes 15,500
Net Cash Provided by Operating Activities 159,900

Net Receipts = Sales + Accounts Receivable (Beginning) - Accounts Receivable (Ending)

Net Receipts = 440,700 + 45,200 - 36,800 = 449,100

Net Payments = Purchase + Inventory (ending) + Accounts Payable (Beginning) - Inventory (Beginning) - Accounts Payable (Ending)

Net Payments = 275,900 + 77,100 + 19,000 - 84,900 - 13,400 = 273,700